OpenStudy (anonymous):

Find the area of a trapezoid in which base1, is 9 units, base2 is 7 units, and the height is 3 units A) 48 units2 <-- Squared B) 24 units2 <-- Squared C) 48 units D) 24 units

OpenStudy (lgbasallote):

i believe the formula is \[\LARGE A = \frac{h(b_1 + b_2)}{2}\] does that help?

OpenStudy (anonymous):

\[A=h(9+7)/2\]

OpenStudy (anonymous):

Like that? then wht do I do?

OpenStudy (anonymous):

16/2 is whatI'm getting & I know its wrong.

OpenStudy (lgbasallote):

A = 3(9+7)/2

OpenStudy (anonymous):

Wait where is the 3 coming from?

OpenStudy (lgbasallote):

height...the h

OpenStudy (anonymous):

Oh nvm, forgot that was there.. Sorry

OpenStudy (lgbasallote):

lol :))

OpenStudy (anonymous):

So A = 3(9+7)/2 3 (16) /2?

OpenStudy (lgbasallote):

yup simplify

OpenStudy (anonymous):

3 (16) /2?

OpenStudy (anonymous):

48/2?

OpenStudy (lgbasallote):

yup..divide

OpenStudy (anonymous):

Wich is 24

OpenStudy (lgbasallote):

there you go

OpenStudy (anonymous):

How can you tell if its squared or not?

OpenStudy (lgbasallote):

\[\LARGE \frac{3 cm \times 16 cm}{2}\] what's cm times cm?

OpenStudy (anonymous):

Oh, also how did you come up with 2 for the bottom of the fraction?

OpenStudy (lgbasallote):

it's the formula for area of trapezoid...or are you asking how i typed that?

OpenStudy (anonymous):

Nope, you answered my question!

OpenStudy (lgbasallote):

glad to help \[\huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]